OpenStudy (anonymous):

Please help! For which value of x does the graph of y = 2x2 − 7x + 6 cross the x-axis? I used the quadratic formula and ended up getting 7 +\- the square root of -97 over 4. I'm stuck in thr rest

OpenStudy (phi):

I used the quadratic formula it looks like you did not get the right answer the quadratic formula is \[ x= \frac{1}{2} \left( -b ± \sqrt{b^2-4ac}\right) \] in your problem \[ 2x^2 − 7x + 6\] the stuff inside the square root is \( (-7)^2 - 4 \cdot 2 \cdot 6= 49 -48= 1 \) and the square root of 1 is 1.